OpenStudy (anonymous):

Find the simplified form of each equation (4/7^5)^2

OpenStudy (anonymous):



OpenStudy (anonymous):

?

OpenStudy (anonymous):

If you're asking if that's the equation then yes. @jose100310

OpenStudy (anonymous):

okay

OpenStudy (ranga):

\[\Large (\frac{ 4 }{ 7^5 })^2 = (\frac{ 2^2 }{ 7^5})^2 = \frac{ 2^4 }{ 7^{10} }\]

OpenStudy (anonymous):

That's not a answer choice? @ranga

OpenStudy (anonymous):

Do you know how to solve it? @jose100310

OpenStudy (ranga):

The problem as typed must be incomplete because it says "equation" but there is no equal sign in the problem!